OpenStudy (hba):

If the radius of a circle is increased by 10 % then the area is increased by,

OpenStudy (tkhunny):

\(\pi r^2\) \(\pi (r*1.1)^{2} = 1.21\cdot\pi r^{2}\) What do you think?

OpenStudy (hba):

@hartnn Yeah lol ? It is a circle but it is a very diff ques..

OpenStudy (hba):

@tkhunny You are welcome to provide hints and explain

OpenStudy (tkhunny):

Already did. Radius increased by 10% -- r*1.1 Area is increased by ???

OpenStudy (hba):

Please explain how did you do the first step

OpenStudy (tkhunny):

Area of a Circle = \(\pi r^{2}\), given the radius. That's the first step.

OpenStudy (hba):

I know that the Area of a circle = pi r^2

OpenStudy (tkhunny):

New Radius is 10% greater than old radius. R = r*1.1 Area of new circle: \(\pi R^{2}\).

OpenStudy (hba):

Got it what next ?

OpenStudy (tkhunny):

Raio of New Area to Old Area \(\dfrac{\pi R^{2}}{\pi r^{2}}\).

OpenStudy (hba):

Ratio*

OpenStudy (tkhunny):

*Ratio - right A little algebra \(\dfrac{\pi R^{2}}{\pi r^{2}} = \dfrac{R^{2}}{r^{2}} = \dfrac{(1.1\cdot r)^{2}}{r^{2}} = \dfrac{1.1^{2}r^{2}}{r^{2}} = 1.1^{2} = 1.21\) And we see a 21% increase!

OpenStudy (hba):

Thank you.